Two parallel straight wires 53.0 cm apart carry currents in opposite directions. Current = 10.0 A is out of the page, and /2 = 8.0 A is into the page. Determine the magnitude and direction of the magnetic field halfway between the two wires.

1
Expert's answer
2022-01-28T12:50:57-0500

Given:

"I_1=10\:\rm A"

"I_2=8.0\:\rm A"

"d=0.53\:\rm m"



The net magnetic field

"B=B_1+B_2"

"B=\frac{\mu_0I_1}{2\pi\ d/2}+\frac{\mu_0I_2}{2\pi d/2}"

"B=\frac{4\pi*10^{-7}*10}{2\pi*0.53/2}+\frac{4\pi*10^{-7}*8.0}{2\pi*0.53/2}=1.4*10^{-5}\:\rm T"

Net magnetic field is directed upward.